OpenStudy (anonymous):

The sum of measure of two angles is 90 degrees. One angle is 28 degrees larger than the other. What would be the measure of each angle?

OpenStudy (anonymous):

x + x + 28 = 90 Smaller = 31 Larger = 59

OpenStudy (anonymous):

31 59

OpenStudy (lalaly):

x+x+28 =90 2x+28=90 x=31 other angle is 31+28 = 59

OpenStudy (anonymous):

x+x+28=90 2x=90-28 2x=62 x=31 31+28=59